OpenStudy (anonymous):

check my work please? question + answer in the comments medal + fan i have a feeling i did something wrong but im not sure what....

jimthompson5910 (jim_thompson5910):

the line you wrote "3456-384 = 3072 cubic inches" is incorrect you don't subtract

OpenStudy (anonymous):

do I divide then?

jimthompson5910 (jim_thompson5910):

"how many times bigger is B compared to A?" you would divide

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so 3456/384 = 9?

OpenStudy (anonymous):

then 9/384 = 0.0234375 round that and its 0.02 so 0.02 * 100 = 2 so it would be 2% ?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hello there

OpenStudy (anonymous):

if u need any help just tag me

jimthompson5910 (jim_thompson5910):

you got 9, so that's equivalent to 900%

OpenStudy (anonymous):

how many more of these do u got

OpenStudy (anonymous):

@Houdini_Dragon thanks for now this is the only question i am stuck on but thank you. :]

OpenStudy (anonymous):

@jim_thompson5910 oh so i dont have to do anything else after receiving the 9?

jimthompson5910 (jim_thompson5910):

just multiply the 9 by 100 to get 900% so not that much in terms of steps after that

OpenStudy (anonymous):

ohh okay thank you so much!! :]